OpenStudy (thadds2003):

Evaluate. 2 – 4 • 3 + 6 • 3 A. 12 B. –28 C. 8 D. 0

OpenStudy (thadds2003):

Will fan and medal. :)

OpenStudy (welshfella):

do the multiplications first 4 . 3 and 6 . 3

OpenStudy (thadds2003):

ok, it equals 216?

OpenStudy (thadds2003):

@welshfella

OpenStudy (thadds2003):

oh man this sucks.

OpenStudy (thadds2003):

@daisyduck04

pooja195 (pooja195):

Use PEMDAS there are no expoenets or paretheses so multiply 6 x 3 =?

OpenStudy (thadds2003):

18

pooja195 (pooja195):

-4 x 3=?

OpenStudy (thadds2003):

-12?

pooja195 (pooja195):

good 2−12+18 2-12=?

OpenStudy (thadds2003):

whoa what?

OpenStudy (thadds2003):

OH 17.

pooja195 (pooja195):

We are just starting form the left

OpenStudy (thadds2003):

-11

pooja195 (pooja195):

2−(4)(3)+(6)(3) =2−12+(6)(3) =−10+(6)(3) =−10+18

Directrix (directrix):

2 – 4 • 3+ 6 • 3 = 2 - (4 * 3) + ( 6 * 3) = 2 - 12 + 18 = -10 + 18 = 8

OpenStudy (thadds2003):

thx

pooja195 (pooja195):

Or just give the answer...

OpenStudy (thadds2003):

ikr, lol

Directrix (directrix):

>>Use PEMDAS That is confusing. Note that I presented a solution, not just an answer.

OpenStudy (thadds2003):

yeah that is true..... :)

Directrix (directrix):

And, an example by which the OP can solve similar problems.

pooja195 (pooja195):

a note for future reference as a fellow helper: sometimes it's best to simply give an explanation and let the person try to find the answer themselves, otherwise they could simply glance at your comment, find the answer they were looking for, thank you, and not even read the explanation you worked so hard to type out :)

OpenStudy (thadds2003):

can i call you brains????